The combustion of 2.034 g of propanol ( C 3 H 7 OH ) increases the temperature of a bomb calorimeter from 298.00 K to 302.88 K . The heat capacity of the bomb calorimeter is 13.99 kJ/K . Determine Δ H for the combustion of propanol to carbon dioxide gas and liquid water.
Answer #1

use:

Q cal = Ccal*delta T

Q cal = 13.99*(302.88-298.0)

Q cal = 68.27 KJ

This heat is supplied by C3H7OH

Molar mass of C3H7OH,

MM = 3*MM(C) + 8*MM(H) + 1*MM(O)

= 3*12.01 + 8*1.008 + 1*16.0

= 60.094 g/mol

mass(C3H7OH)= 2.034 g

use:

number of mol of C3H7OH,

n = mass of C3H7OH/molar mass of C3H7OH

=(2.034 g)/(60.09 g/mol)

= 3.385*10^-2 mol

use:

delta H = -Qcal/number of mol

= -68.27 KJ / 3.385*10^-2 mol

= -2.017*10^3 KJ/mol

Answer: -2.02*10^3 KJ/mol